y+5x=0
x+y=12
how do you do this? please show all steps

Found 2 solutions by umnikmo, oberobic:
Answer by umnikmo(75)

from first eq: y=-5x
substitute this into second eq:
x+(-5x)=12
-4x=12 => x=-3
y=-5*x=-5*(-3)=15

Answer by oberobic(2304)

y + 5x = 0
x + y = 12
.
re-arrange the first equation
and subtract the second equation
.
5x +y = 0
x + y = 12
-----------
4x = -12
.
divide by 4
.
x = -3
.
substitute
.
-3 +y = 12
.
y = 15
.
Check these values using the other equation.
.
15 + 5*(-3) = 0
True
.
Answer: The values are x = -3 and y = 15.
.
Done.